# Alignment without Over-optimization: Training-Free Solution for Diffusion Models

### v1

@@ A.3 Pseudocode for Online black-box optimization
For both aesthetic score experiments (Section 4.3 and JPEG compressibility experiments (SectionF.3), we used MLP on frozen CLIP embeddings (Radford et al., 2021) for surrogate model.
@@ C.2.2 Assumptions for Proposition 2 and 3
We go over the viability of these assumptions. Assumption (a) can be satisfied lower and upper bounded rewards by adding a constant. Real-world rewards are indeed lower and upper bounded in most practical settings, including aesthetic score and PickScore used in our experiments. Even if not, we can simply clamp the reward to ensure the condition. Assumption (b) should be ensured for numerical stability of the algorithm. That is, if the gradient explode, generation using the guidance isn’t possible. Since $r(\cdot)$ and $\hat{x}(\cdot)$ are function using neural networks, the assumption can be met unless $x_{t}$ gets out of support of the training distribution. Experimentally, this is commonly true unless using extremely small $\alpha$. Note that the assumption of uniform bound in time is only for simplicity and the bound may change in time. Assumption (c) is generally not true since $m_{t}$ is a Gaussian kernel and $p_{t}$ is also the marginal distribution over Gaussian noise added to clean data. However, it can ’effectively’ be satisfied. We explain what this means in more detail. First, the data manifold $\mathcal{X}_{0}$ can be assumed to be bounded since most real-world data without corruption doesn’t contain infinitely large or small component. Next, suppose we define the forward, reverse diffusion process and the Markov kernels using Gaussian distribution truncated at tail probability $\epsilon$ instead of standard Gaussian. Numerically, when $\epsilon$ is sufficiently small, the impact of the truncation becomes negligible, hence the diffusion process and the SMC sampler behaves similarly to the original Gaussian case. However, unlike the unbounded Gaussian noise, the bounded support of the truncated noise ensures compactness $\mathcal{X}_{t}$ in Assumption (c). Thus the models can be modified to satisfy Assumption (c) without any effect of the practical algorithm.
@@ Appendix D Implementation Details
Appendix E Comparison with Diffusion-based Samplers
@@
Experiment Setting. We conducted additional experiments using GMM examples from Figure 1. To check our hypothesis and demystify the effect of each element, we conducted experiment using fine-tuning + NN parameterization (FT-NN, equivalent to direct backpropagation with KL regularization), fine-tuning + Grad parameterization (FT-Grad), training from scratch + NN parameterization (Train-NN), training from scratch + Grad parameterization (Train-Grad, equivalent to DDS). We excluded PIS from our comparison since DDS already outperforms PIS, where PIS uses pinned Brownian motion running backward in time as reference diffusion, which is proven to incur instability both theoretically (Appendix A.2 of Vargas et al. (2023)) and empirically (Appendix C.4 of Vargas et al. (2023)).
@@
Appendix F Additional Experiment Results
F.1 3d Swiss Roll
@@ F.3 Non-differentiable Rewards
F.4 Aesthetic Score additional results
@@
F.5 PickScore additional results
@@
Appendix G Future works

### v2

@@ A.3 Pseudocode for Online black-box optimization
For both aesthetic score experiments (Section 4.3 and JPEG compressibility experiments (SectionG.3), we used MLP on frozen CLIP embeddings (Radford et al., 2021) for surrogate model.
@@ C.2.2 Assumptions for Proposition 2 and 3
We go over the viability of these assumptions. Assumption (a) can be satisfied lower and upper bounded rewards by adding a constant. Real-world rewards are indeed lower and upper bounded in most practical settings, including aesthetic score and PickScore used in our experiments. Even if not, we can simply clamp the reward to ensure the condition. Assumption (b) should be ensured for numerical stability of the algorithm. That is, if the gradient explode, generation using the guidance isn’t possible. Since $r(\cdot)$ and $\hat{x}(\cdot)$ are function using neural networks, the assumption can be met unless $x_{t}$ gets out of support of the training distribution. Experimentally, this is commonly true unless using extremely small $\alpha$. Note that the assumption of uniform bound in time is only for simplicity and the bound may change in time. Assumption (c) is generally not true since $m_{t}$ is a Gaussian kernel and $p_{t}$ is also the marginal distribution over Gaussian noise added to clean data. However, it can ’effectively’ be satisfied. We explain what this means in more detail. First, the data manifold $\mathcal{X}_{0}$ can be assumed to be bounded since most real-world data without corruption doesn’t contain infinitely large or small component. Next, suppose we define the forward, reverse diffusion process and the Markov kernels using Gaussian distribution truncated at tail probability $\epsilon$ instead of standard Gaussian. Numerically, when $\epsilon$ is sufficiently small, the impact of the truncation becomes negligible, hence the diffusion process and the SMC sampler behaves similarly to the original Gaussian case. However, unlike the unbounded Gaussian noise, the bounded support of the truncated noise ensures compactness of $\mathcal{X}_{t}$ in Assumption (c). Thus the algorithm can be modified to satisfy Assumption (c) without any effect of the practical algorithm.
@@ Appendix D Implementation Details
Appendix E Test-time Scaling
Test-time Scaling. While DAS offers a powerful test-time reward maximization scheme, even outperforming fine-tuning methods, it requires additional computation during inference. Similarly, recent literature on language models suggests that leveraging additional computation during inference can further improve the response’s quality (Wei et al., 2022; Khanov et al., 2024; Wu et al., 2024; Snell et al., 2024). However, it’s important to verify which method can most effectively utilize additional resources and achieve the highest performance as we scale test-time computation.
Experiment Setup. To demonstrate DAS is indeed an effective way to leverage computation during inference, we compare the test-time scaling of DAS with other possible test-time alignment methods for diffusion models. Baseline methods include Best-of-N (Touvron et al., 2023; Beirami et al., 2024; Ma et al., 2025) which runs $N$ (number of particles) independent generation process and selects the final sample with the highest reward. SMC uses a Sequential Monte Carlo sampler as DAS does, but without the approximate locally optimal proposal in Section 3.3.3 and tempering, using the original generation process as the proposal for each step. Including DAS, these methods are naturally scalable by increasing the number of particles used in the algorithm. We use the number of particles in $\{1,2,4,8,16,32,64,128\}$ for each method and compare how target reward and unseen reward for cross-reward generalization scale as inference-time computation increase. We use PickScore for target reward and HPSv2 for unseen reward.
Results. Figure 8 shows that as the inference compute increases, DAS consistently outperforms both Best-of-N and SMC across both PickScore (target reward) and HPSv2 (unseen reward). Notably, DAS already achieves superior performance over other baselines even when using fewer particles and significantly less compute. This highlights DAS’s efficiency in maximizing rewards compared to methods like Best-of-N, which shows limited improvements beyond a certain compute threshold, and SMC, which scales better than Best-of-N but still falls short of DAS.
These results emphasize the strength of the proposal and tempering scheme utilized by DAS, which enables it to effectively improve sample efficiency and thus leverage test-time computation without requiring extensive resources. By achieving better performance at lower computational cost and scaling seamlessly with increased resources, DAS proves to be both a compute-efficient and scalable method for inference-time alignment in diffusion models.
Figure 8: DAS outperforms with fewer inference-time compute. We compare DAS with other inference-time alignment algorithms while changing the inference-time compute, units in Teraflops. (left) Inference-time Compute vs. PickScore (target reward), (right) Inference-time Compute vs. HPSv2 (unseen reward)
Appendix F Comparison with Diffusion-based Samplers
@@
Experiment Setup. We conducted additional experiments using GMM examples from Figure 1. To check our hypothesis and demystify the effect of each element, we conducted experiment using fine-tuning + NN parameterization (FT-NN, equivalent to direct backpropagation with KL regularization), fine-tuning + Grad parameterization (FT-Grad), training from scratch + NN parameterization (Train-NN), training from scratch + Grad parameterization (Train-Grad, equivalent to DDS). We excluded PIS from our comparison since DDS already outperforms PIS, where PIS uses pinned Brownian motion running backward in time as reference diffusion, which is proven to incur instability both theoretically (Appendix A.2 of Vargas et al. (2023)) and empirically (Appendix C.4 of Vargas et al. (2023)).
@@
Appendix G Additional Experiment Results
G.1 3d Swiss Roll
@@ G.3 Non-differentiable Rewards
G.4 Aesthetic Score additional results
@@
G.5 PickScore additional results
@@
Appendix H Future works